Circular No. 6420

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) K. Jockers and G. Lichtenberg, Max Planck Institute for Aeronomy,

Katlenburg-Lindau, write: “From observations of this comet obtained on Apr. 1.9 UT with the Hoher List 1-m reflector, we derive a CO^{+}/H_{2}O^{+} ratio > 1 at a tailward distance of 6.4 × 10^{4} km from the nucleus. In comet 1P/Halley, this value was < 0.2 (Lutz et al. 1993, Ap.J. 407, 402). This indicates that C/1996 B2 may be rich in carbon compounds.”

Visual m_{1} estimates: June 9.79 UT, 5.3 (D. Seargent, The Entrance, N.S.W., Australia, naked eye); 11.80, 5.2 (P. Camilleri, Cobram, Vic., Australia, 20×80 binoculars).

ROSAT ALL-SKY SURVEY BRIGHT SOURCE CATALOGUE W. Voges, B. Aschenbach, T. Boller, H. Bräuninger, U. Briel, W.

Burkert, K. Dennerl, J. Englhauser, R. Gruber, F. Haberl, G. Hartner, G. Hasinger, M. Kürster, E. Pfeffermann, W. Pietsch, P. Predehl, C. Rosso, J. H. M. M. Schmitt, J. Trümper, and H.-U. Zimmermann, Max-Planck-Institut für Extraterrestrische Physik, Garching, report: “The ROSAT All-Sky Survey Bright Source Catalogue (RASS-BSC, revision 1RXS) has been released and is available through the World Wide Web (URL http:

//www.rosat.mpe-garching.mpg.de/survey/rass-bsc/) and via anony-
mous ftp (host ftp.rosat.mpe-garching.mpg.de, directory archive/

survey/rass-bsc). This catalogue is derived from the all-sky survey performed during the first half year of the ROSAT mission in 1990–1991; 18 811 sources are catalogued, with a limiting ROSAT PSPC countrate of 0.05 cts/s in the energy band 0.1–2.4 keV. The sources have a detection likelihood of at least 15 and contain at least 15 source photons. At a brightness limit of 0.1 cts/s (8547 sources) the catalogue represents a sky coverage of 92 percent. For each source are provided the ROSAT name, the position in equatorial coordinates, the positional error, the source countrate and error, the background countrate, exposure time, hardness-ratios HR1 and HR2 and errors, extent and likelihood of extent, and likelihood of detection. For 94 percent of the sources, visual inspection confirmed the results of the standard processing with respect to existence and position; the remaining 6 percent were reanalyzed and appropriately flagged. Broadband images are available for a subset of the flagged sources. A list containing correlations with other catalogues and identifications will follow soon. Questions

or comments may be directed to survey@rosat.mpe-garching.mpg.de.”

1996 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green